Calculate the pH for each case in the titration of 50.0 mL of 0.140 M HClO(aq) with 0.140 M KOH(aq). Use the ionization constant for HClO.

What is the pH before addition of any KOH?

pH=

What is the pH after addition of 25.0 mL KOH?

pH=

What is the pH after addition of 40.0 mL KOH?

pH=

What is the pH after addition of 50.0 mL KOH?

pH=

What is the pH after addition of 60.0 mL KOH?

pH=

Please show work
Answer #1

Helo: Ka= 4x108 pla= -log ka = -log (4x158)= 7.398 © Helot Ho - Hot+ clo 0 1 +x 0.14 Hub XX E 0.14-X EH, ot] [c10] {helo] o.1pH s pria tiog mellom = 7.38 tleg en pH = 7.398 Kod: Molarity = 0.14 M volume= 40mia 8.042 no-ot moles = 0.14X0-04=0.0056 HelKOH: inolarity = 0.14M CU volume = 50mla 0-052 no ob molesa o.lux 8.05= 0.007 - Itclot Koh - kclot to weak acid strong base s- Kolt: molarity = 0.14M volumea fomla 0.062 no-ct moles a Olux 0-06 = 0.0084 Helot Kot - kolot H₂O 0-007 10.0084 C-000 -0-00pH: formulas for salts weak acidt strong base ->salft ito pH = Ź [pfw tpka + log c] strong acid & weakbase salf to pH = ₃ [